OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = x + 4 and passes through the point (-6, 3) ? help

OpenStudy (anonymous):

first you need the slope of the perpendicular line slope*slope_p=-1 the slope of the eq. is 1 then slope of perpendicular is -1

OpenStudy (anonymous):

the eq. y-y1=m(x-x1) y-(3)=-1(x-(-6)) y-3=-x-6 y=-x-3